OpenStudy (anonymous):

find the exact value of cos 15 degrees

OpenStudy (anonymous):

use a cos difference identity

OpenStudy (anonymous):

Cos(45 degrees)-Cos(30 degrees)

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

no cos(45-30)

OpenStudy (anonymous):

That's what i meant. Sorry!

OpenStudy (anonymous):

do you knwo your cos difference identity?

OpenStudy (anonymous):

And do you know your unique angles? (30, 45, 60, and 90)?

OpenStudy (anonymous):

\[\cos(A-B)=CosACosB+SinASinB \]

OpenStudy (anonymous):

wait. i thought you were the op lol. i guess we should wait for the op

OpenStudy (anonymous):

Heheh.

OpenStudy (anonymous):

waitt.... i know how to do this...

OpenStudy (anonymous):

i hope so.

OpenStudy (anonymous):

CALCULATOR

OpenStudy (anonymous):

that is a complete disservice to your education.

OpenStudy (anonymous):

LOL...u never mentioned which method to use...this is just me using my intelligence

OpenStudy (anonymous):

because it isnt my question. i know how to solve it. they will find calculus much easier if they can get identies down.

OpenStudy (anonymous):

Why is this still open?

OpenStudy (anonymous):

\[ \cos^2(x) = \frac{1+\cos(2x)}{2} \]In our case: \[ \cos^2(15) = \frac{1+\cos(30)}{2} \]

OpenStudy (anonymous):

This means \[ |\cos^2(15)| = \sqrt{\frac{1+\cos(30)}{2}} \]We can tell that \(15\) is in the first quadrant, so our answer should be positive.

OpenStudy (anonymous):

 For a 30 degree right angle, first imagine an equilateral triangle with 60 degrees in each angle.

OpenStudy (anonymous):

 Cut it in half, and realize that you have two 30 degree angles.

OpenStudy (anonymous):

It's clear that the hypotenuse is twice the length of the opposite side. So since \[ hypotenuse^2 = adjacent^2+opposite^2 \]We can say: \[ hypotenuse^2 = adjacent^2 + \left(\frac{hypotenuse}{2}\right)^2 \]

OpenStudy (anonymous):

Remember that \(\cos(x) = \frac{adjacent}{hypotenuse}\). If we let \(hypotenuse = 1\), then \(\cos(x) = adjacent\).

OpenStudy (anonymous):

So \[ 1^2 = adjacent^2 + \left(\frac{1}{2}\right)^2 \]Or \[ adjacent^2 = 1 - \frac{1}{4} \]Solving for the adjacent side, we get: \[ adjacent = \sqrt{\frac{3}{4}} = \frac{\sqrt 3}{2} \]

OpenStudy (anonymous):

In short: \(\cos(30^\circ) = \frac{\sqrt 3}{2}\)